sum of the digits of a two no. is 9 the no. obtained by interchanging the digits exceed the given no. by 27 find the given no.
Answer by checkley71(8403)

X+Y=9 OR X=9-Y
10X+Y=10Y+X+27
10(9-Y)+Y=10Y+(9-Y)+27
90-10Y+Y=10Y+9-Y+27
-18Y=36-90
-18Y=-54
Y=-54/-18
Y=3 ANSWER.
X+3=9
X=9-3
X=6 ANSWER.
PROOF
10*6+3=10*3+6+27
60+3=30+6+27
63=63